# No exponential quantum speedup for $\mathrm{SIS}^{\infty}$ anymore

### v2

@@ 1 Introduction
In 2021, Chen, Liu, and Zhandry [CLZ22] did precisely this. They presented an efficient quantum algorithm for the well-known Short Integer Solution in $\ell_{\infty}$ norm ($\mathrm{SIS}^{\infty}$) problem, in a less-studied parameter regime, but one with no known classical polynomial-time algorithm. Here is an example result they proved (see Theorem 1.4 for more). Let $q$ be a prime. [^1] We use $q$ (instead of $p$) to be consistent with cryptography literature. Our results also work for prime power or composite modulus. See Section 5 for details.
@@
The particularly exciting part about this work is that it used genuinely different techniques from existing quantum speedups. Furthermore, in a more standard and more challenging parameter regime — e.g., $m=Cn$ and “short” meaning [^3] Here $x$’s entries are interpreted in $\{-\lfloor q/2\rfloor,\dots,\lfloor q/2\rfloor\}$. $\|x\|_{\infty}\leq q/4$ — the $\mathrm{SIS}^{\infty}$ problem is of significant cryptographic interest. Variants of the problem underlie the security of several (candidate post-quantum) cryptographic systems, such as Dilithium [DKL+18] and Wave [DST19].
The Chen–Liu–Zhandry quantum algorithm is based on a quantum reduction by Regev [Reg09], who based the hardness of the Learning With Errors problem on the hardness of finding short lattice vectors. A similar reduction shows that $\mathrm{SIS}^{\infty}$ can be solved efficiently if a certain decoding problem can be solved efficiently. Instead of using Regev’s reduction to show hardness, Chen, Liu, and Zhandry devise a clever, efficient quantum algorithm for the decoding problem (in a certain parameter regime), and thereby obtain an efficient quantum algorithm for $\mathrm{SIS}^{\infty}$.
These algorithmic applications of Regev’s reduction have since attracted considerable interest [CLZ22, CT23, YZ24, JSW+24, CT24]. Besides the algorithm for $\mathrm{SIS}^{\infty}$, two other quantum algorithms that use the same primitive are: (1) the algorithm of Yamakawa and Zhandry [YZ24] that achieves an exponential quantum–classical black-box separation for a search problem relative to a random oracle; and, (2) the Decoded Quantum Interferometry (DQI) algorithm for various optimization problems, such as the Optimal Polynomial Intersection (OPI) problem [JSW+24, CT24]. The result by Yamakawa and Zhandry can be made non-oracular by instantiating the random oracle with a cryptographic hash function, providing a concrete problem that presumably retains the exponential quantum speedup. This leaves us with at least three non-oracular problems based on this algorithmic primitive with potential exponential speedups: $\mathrm{SIS}^{\infty}$, OPI, and Yamakawa–Zhandry.
Our main result is to dequantize [^4] We do not simulate the CLZ algorithm in a classical way. Rather, we provide a classical algorithm directly. [CLZ22]. We show that $\mathrm{SIS}^{\infty}$, and a further generalization they study, can be solved by a classical algorithm whose efficiency even outperforms that of their quantum algorithm. For example, we prove (see Theorem 1.5 for more) the following theorem.
@@
We highlight several advantages of our result: (1) it is classical and deterministic; (2) it works for worst-case $H$; (3) the requirement on $m$ has no dependence on $q$; (4) it is $\operatorname{poly}(n)$-time even when $q=2^{\operatorname{poly}(n)}$; and (5) the notion of “short” is much stricter. (Our algorithm’s classical running time is even faster than CLZ’s quantum running time in the setting of comparable $m$; see Section 5.) As we also discuss in Section 5, for this particular sample theorem, a similar but slightly weaker version (without features (3)(4)) can be extracted from a recent work of Imran and Ivanyos [II24].
In general, the main theorem in this work gives a classical algorithm that outperforms the most general quantum algorithm from [CLZ22]. This quantum algorithm is for a generalization of $\mathrm{SIS}^{\infty}$ we call Constrained Integer Solution ($\mathrm{CIS}$) problem, in which the entries of $x$ are restricted to lie in a general set $A$, as opposed to an interval. (For more details, see Section 1.4.)
@@
([CLZ22]). Let $k\geq 2$ be a constant and $A\subseteq\mathbb{F}_{q}$ be of size |A|=qk+1|A|=q-k+1. There is a $\operatorname{poly}(m,q)$-time quantum algorithm that, given a uniformly random $H\in\mathbb{F}_{q}^{n\times m}$ where
@@
Theorem 1.1 (Special case of [CLZ22, Remark 4]).
@@ 1.2 Short Integer Solution in $\ell_{\infty}$ norm, $\mathrm{SIS}^{\infty}$
This problem is interesting both in the worst case and the average case. [^7] Assuming worst-case hardness of some lattice problems and passing through the SIS problem in $\ell_{2}$ norm [Ajt96, MR07], the average-case $\mathrm{SIS}^{\infty}$ problem is hard when $s=O(1)$, $q=\Theta(n^{2}\log n)$, and $m=\Theta(n\log n)$. The main result of Chen, Liu, and Zhandry is a quantum algorithm for the problem in the average case, where the matrix $H$ is drawn uniformly at random from $\mathbb{F}_{q}^{n\times m}$.
Theorem 1.4 ([CLZ22, Theorem 2]).
@@
We present the proof of this weaker theorem in Section 3, as a warm-up to the more elaborate proof needed for Theorem 1.5. As we discuss in Section 5, this easier proof is based on a method of Imran and Ivanyos [II24], but improves upon it by allowing exponentially large $q$.
@@
There are two prior algorithms we can compare against: a quantum algorithm by Chen, Liu, and Zhandry [CLZ22] and a classical one by Imran and Ivanyos [II24].
Theorem 1.7 (Follows from [CLZ22, Remark 4]).
@@
Theorem 1.8 (Follows from [II24, Remark 4]).
@@ 1.4 Constrained Integer Solution ($\mathrm{CIS}$)
While the main result presented by Chen, Liu, and Zhandry is about $\mathrm{SIS}^{\infty}$, their algorithm works for a more general problem, as pointed out in [CLZ22, Remark 4]. In this version, instead of requiring that the solution $x$ has small infinity norm, we require that each entry of $x$ belongs to a specified allowed set $A\subseteq\mathbb{F}_{q}$. We call this problem “Constrained Integer Solution ($\mathrm{CIS}$)” or sometimes $A$-$\mathrm{CIS}$ to emphasize the set $A$.
@@
Like [CLZ22], we will only study this problem in the average case, where $H$ is a uniformly random matrix in $\mathbb{F}_{q}^{n\times m}$, as the worst-case instances may not guarantee solutions. We also remark that the problem seems quite hard unless $|A|$ is close to $q$; thus we parameterize |A|=qk+1|A|=q-k+1.
It turns out that the same quantum algorithm in [CLZ22] that solves $\mathrm{SIS}^{\infty}$ can be used for $\mathrm{CIS}$ with similar bounds.
Theorem 1.11 ([CLZ22, Remark 4]).
@@
We give classical algorithms for the same problem. First, we have a simple argument to handle the case when $q$ is much larger than $k$. This is in fact the motivating setting in [CLZ22], as they typically describe $q$ being polynomial in $n$ and $k$ being constant.
@@ 1.5 Results summary
One way to improve Theorem 2.1 is to observe a dimension reduction: once $u_{1}$ is created, we can project later vectors onto the space orthogonal to $u_{1}$; this is because $0\cdot u_{1},1\cdot u_{1},-1\cdot u_{1}$ can all be replaced by a subset-sum from $S_{1}$. This leads to the following Theorem 2.2, which is a special case of [ISS07, Claim 1] and we present a short proof here.
@@ 3 A halving trick and its applications
The goal of this section is to give a primitive exposition of the halving trick using reducible vectors. We note that our halving trick is inspired by a similar one in [II24], and our main contribution here is to remove the dependence on the field size $q$ in [II24].
@@
$c\cdot u=c\cdot u-\sum_{i\in[m]}\alpha_{i}v_{i}=\sum_{i\colon|\alpha_{i}|\leq h^{\prime}}(-\alpha_{i})v_{i}+\sum_{i\colon\alpha_{i}>h^{\prime}}(c-\alpha_{i})v_{i}+\sum_{i\colon\alpha_{i}<h^{\prime}}-(c+\alpha_{i})v_{i},$
@@
Assume $u^{(1)},\ldots,u^{(m)}$ are all nonzero vectors. Then $\mathcal{B}$ applies $\mathcal{A}$ again, to $u^{(1)},\dots,u^{(m)}$, thereby getting a nontrivial $(\pm h)$-zero-sum $\sum_{i}c_{i}u^{(i)}=0$, where the $c_{i}$’s are not all zero. Since each $u^{(i)}$ is $(\pm h\to\pm\lfloor h/2\rfloor)$-reducible, we can replace $c_{i}u^{(i)}$ with a $(\pm\lfloor h/2\rfloor)$-sum of vectors from $S_{i}$. After the substitution, we obtain a $(\pm\lfloor h/2\rfloor$-zero-sum. To see it is nontrivial, we recall that some $c_{i}$ is nonzero and the corresponding $u^{(i)}$ is also nonzero, which means $c_{i}u^{(i)}$ is nonzero and hence its substitution cannot be a trivial sum.
@@ 3.1 Simple $\mathrm{SIS}^{\infty}$ algorithms
We also remark that Theorem 3.9 (and even its weaker version Theorem 3.5) improves the result of [II24] by a factor of $q^{O(q\log q)}$.
@@
$\sum_{i\in S}u_{i}^{\prime}-\sum_{i\in T}u_{i}^{\prime}=0,$
where $S,T\subseteq[m]$ are disjoint and not both empty. With Equation 11, we obtain the following zero-sum
$\sum_{i\in S}u_{i}^{\prime}+\sum_{i\in T}u_{i}=\left(\sum_{i\in S}u_{i}^{\prime}-\sum_{i\in T}u_{i}^{\prime}\right)+\sum_{i\in T}(u_{i}+2\cdot u_{i}^{\prime})=0,$
@@
This means $\sum_{i\in S_{j}}\alpha_{i}v_{i}=c_{j}v^{*}$ where $\beta_{j}=\sum_{i\in S_{j}}\alpha_{i}c_{i}$. Suppose there is some set $T\subseteq[d]$ that $\sum_{j\in T}\beta_{j}=1$. We can use the textbook dynamic programming algorithm to find such a set in time $\operatorname{poly}(q,d)$, then obtain a desired $(\pm 1)$-sum
@@
The reduction in [II24] produces a bound of $m\geq q^{O(q\log^{2}q)}n^{O(q\log q)}$, whereas our Theorem 3.12 produces $m\geq n^{O(q)}$.
@@ 3.3 Simple $\mathrm{CIS}$ algorithms: large primes
Let $q\geq 3$ be a prime and let $A\subseteq\mathbb{F}_{q}$ be nonempty. Suppose there is a deterministic algorithm $\mathcal{A}$, given $m$ input vectors in $\mathbb{F}_{q}^{n}$ and running in time $T$, that outputs a nontrivial $A$-zero-sum.
@@
Without loss of generality assume $A$ is the $(2k+1)$-AP. By Lemma 3.13, it suffices to have a worst-case algorithm for nontrivial $(\pm\lfloor q/(2k)\rfloor)$-zero-sums, for which we use Theorem 3.9. ∎
In the case of large $q$, we have known results in arithmetic combinatorics that guarantee long APs. The following fact is due to Lev [Lev00] and, in Appendix A, we provide a short proof for completeness.
Fact 3.16 ([Lev00]).
@@ 4.1 Basic zero-sum algorithms for specific nontriviality
Denote $A=\{a_{1},\ldots,a_{s}\}$ where $s=|A|$. Let $\ell\leq n$ be the rank of the input vectors. By an invertible linear transform and in $\operatorname{poly}(n,r,|A|)$ time, we can assume that we are given standard basis $e_{1},\ldots,e_{\ell}$ and $v_{1}^{(i)},\ldots,v_{s}^{(i)},i\in[r]$ as $rs$ extra vectors that lie in the span of $e_{1},\ldots,e_{\ell}$.
@@
Since $\beta_{1},\ldots,\beta_{r}$ are not all zero, we can safely assume $\beta_{j}=1$ for some $j\in[r]$. Hence by substituting Equation 12 into Equation 13, we know that $a_{1},\ldots,a_{s}$ all appear as coefficients of $u_{\beta}$, which implies that $u_{\beta}$ is $A$-nontrivial.

### v3

@@ 1 Introduction
In 2021, Chen, Liu, and Zhandry [9] did precisely this. They presented an efficient quantum algorithm for the well-known Short Integer Solution in $\ell_{\infty}$ norm ($\mathrm{SIS}^{\infty}$) problem, in a less-studied parameter regime, but one with no known classical polynomial-time algorithm. Here is an example result they proved (see Theorem 1.4 for more). Let $q$ be a prime. [^1] We use $q$ (instead of $p$) to be consistent with cryptography literature. Our results also work for prime power or composite modulus. See Section 5.1 for details.
@@
The particularly exciting part about this work is that it used genuinely different techniques from existing quantum speedups. Furthermore, in a more standard and more challenging parameter regime — e.g., $m=Cn$ and “short” meaning [^3] Here $x$’s entries are interpreted in $\{-\lfloor q/2\rfloor,\dots,\lfloor q/2\rfloor\}$. $\|x\|_{\infty}\leq q/4$ — the $\mathrm{SIS}^{\infty}$ problem is of significant cryptographic interest. Variants of the problem underlie the security of several (candidate post-quantum) cryptographic systems, such as Dilithium [12] and Wave [10]; see Section 5.6 for more details.
The Chen–Liu–Zhandry quantum algorithm is based on a quantum reduction by Regev [30], who based the hardness of the Learning With Errors problem on the hardness of finding short lattice vectors. A similar reduction shows that $\mathrm{SIS}^{\infty}$ can be solved efficiently if a certain decoding problem can be solved efficiently. Instead of using Regev’s reduction to show hardness, Chen, Liu, and Zhandry devise a clever, efficient quantum algorithm for the decoding problem (in a certain parameter regime), and thereby obtain an efficient quantum algorithm for $\mathrm{SIS}^{\infty}$.
These algorithmic applications of Regev’s reduction have since attracted considerable interest [9, 7, 35, 22, 8]. Besides the algorithm for $\mathrm{SIS}^{\infty}$, two other quantum algorithms that use the same primitive are: (1) the algorithm of Yamakawa and Zhandry [35] that achieves an exponential quantum–classical black-box separation for a search problem relative to a random oracle; and, (2) the Decoded Quantum Interferometry (DQI) algorithm for various optimization problems, such as the Optimal Polynomial Intersection (OPI) problem [22, 8]. The result by Yamakawa and Zhandry can be made non-oracular by instantiating the random oracle with a cryptographic hash function, providing a concrete problem that presumably retains the exponential quantum speedup. This leaves us with at least three non-oracular problems based on this algorithmic primitive with potential exponential speedups: $\mathrm{SIS}^{\infty}$, OPI, and Yamakawa–Zhandry.
Our main result is to dequantize [^4] We do not simulate the CLZ algorithm in a classical way. Rather, we provide a classical algorithm directly. [9]. We show that $\mathrm{SIS}^{\infty}$, and a further generalization they study, can be solved by a classical algorithm whose efficiency even outperforms that of their quantum algorithm. For example, we prove (see Theorem 1.5 for more) the following theorem.
@@
We highlight several advantages of our result: (1) it is classical and deterministic; (2) it works for worst-case $H$; (3) the requirement on $m$ has no dependence on $q$; (4) it is $\operatorname{poly}(n)$-time even when $q=2^{\operatorname{poly}(n)}$; and (5) the notion of “short” is much stricter. (Our algorithm’s classical running time is even faster than CLZ’s quantum running time in the setting of comparable $m$; see Section 5.1.) As we also discuss in Section 5.2, for this particular sample theorem, a similar but slightly weaker version (without features (3)(4)) can be extracted from a recent work of Imran and Ivanyos [17].
In general, the main theorem in this work gives a classical algorithm that outperforms the most general quantum algorithm from [9]. This quantum algorithm is for a generalization of $\mathrm{SIS}^{\infty}$ we call Constrained Integer Solution ($\mathrm{CIS}$) problem, in which the entries of $x$ are restricted to lie in a general set $A$, as opposed to an interval. (For more details, see Section 1.4.)
@@
([9]). Let $k\geq 2$ be a constant and $A\subseteq\mathbb{F}_{q}$ be of size |A|=qk+1|A|=q-k+1. There is a $\operatorname{poly}(m,q)$-time quantum algorithm that, given a uniformly random $H\in\mathbb{F}_{q}^{n\times m}$ where
@@
Theorem 1.1 (Special case of [9, Remark 4]).
@@ 1.2 Short Integer Solution in $\ell_{\infty}$ norm, $\mathrm{SIS}^{\infty}$
This problem is interesting both in the worst case and the average case. [^7] Assuming worst-case hardness of some lattice problems and passing through the SIS problem in $\ell_{2}$ norm [1, 27], the average-case $\mathrm{SIS}^{\infty}$ problem is hard when $s=O(1)$, $q=\Theta(n^{2}\log n)$, and $m=\Theta(n\log n)$. The main result of Chen, Liu, and Zhandry is a quantum algorithm for the problem in the average case, where the matrix $H$ is drawn uniformly at random from $\mathbb{F}_{q}^{n\times m}$.
Theorem 1.4 ([9, Theorem 2]).
@@
We present the proof of this weaker theorem in Section 3, as a warm-up to the more elaborate proof needed for Theorem 1.5. As we discuss in Section 5.2, this easier proof is based on a method of Imran and Ivanyos [17], but improves upon it by allowing exponentially large $q$.
@@
There are two prior algorithms we can compare against: a quantum algorithm by Chen, Liu, and Zhandry [9] and a classical one by Imran and Ivanyos [17].
Theorem 1.7 (Follows from [9, Remark 4]).
@@
Theorem 1.8 (Follows from [17, Proposition 4]).
@@ 1.4 Constrained Integer Solution ($\mathrm{CIS}$)
While the main result presented by Chen, Liu, and Zhandry is about $\mathrm{SIS}^{\infty}$, their algorithm works for a more general problem, as pointed out in [9, Remark 4]. In this version, instead of requiring that the solution $x$ has small infinity norm, we require that each entry of $x$ belongs to a specified allowed set $A\subseteq\mathbb{F}_{q}$. [^8] We remark that the results in [9] work for the more general setting where each entry of $x$ gets its own restricted subset. Our algorithms also work in this setting; see Section 5.1 for a discussion. We call this problem “Constrained Integer Solution ($\mathrm{CIS}$)” or sometimes $A$-$\mathrm{CIS}$ to emphasize the set $A$.
@@
Like [9], we will only study this problem in the average case, where $H$ is a uniformly random matrix in $\mathbb{F}_{q}^{n\times m}$, as the worst-case instances may not guarantee solutions. We also remark that the problem seems quite hard unless $|A|$ is close to $q$; thus we parameterize |A|=qk+1|A|=q-k+1.
It turns out that the same quantum algorithm in [9] that solves $\mathrm{SIS}^{\infty}$ can be used for $\mathrm{CIS}$ with similar bounds.
Theorem 1.11 ([9, Remark 4]).
@@
We give classical algorithms for the same problem. First, we have a simple argument to handle the case when $q$ is much larger than $k$. This is in fact the motivating setting in [9], as they typically describe $q$ being polynomial in $n$ and $k$ being constant.
@@ 1.5 Results summary
One way to improve Theorem 2.1 is to observe a dimension reduction: once $u_{1}$ is created, we can project later vectors onto the space orthogonal to $u_{1}$; this is because $0\cdot u_{1},1\cdot u_{1},-1\cdot u_{1}$ can all be replaced by a subset-sum from $S_{1}$. This leads to the following Theorem 2.2, which is a special case of [19, Claim 1] and we present a short proof here.
@@
We remark that the above lemma (and the upcoming generalization Lemma 3.6) can be viewed as a coding-theoretic result to efficiently find a sparse codeword. Along this line, similar result can be derived from algorithmizing the Plotkin bound [29], though they are stuck at the same $2/3$ ratio.
@@ 3 A halving trick and its applications
The goal of this section is to give a primitive exposition of the halving trick using reducible vectors. We note that our halving trick is inspired by a similar one in [17], and our main contribution here is to remove the dependence on the field size $q$ in [17].
@@
$c\cdot u=c\cdot u-\sum_{i\in[m]}\alpha_{i}v_{i}=\sum_{i\colon|\alpha_{i}|\leq h^{\prime}}(-\alpha_{i})v_{i}+\sum_{i\colon\alpha_{i}>h^{\prime}}(c-\alpha_{i})v_{i}+\sum_{i\colon\alpha_{i}<-h^{\prime}}-(c+\alpha_{i})v_{i},$
@@
Assume $u^{(1)},\ldots,u^{(m)}$ are all nonzero vectors. Then $\mathcal{B}$ applies $\mathcal{A}$ again, to $u^{(1)},\dots,u^{(m)}$, thereby getting a nontrivial $(\pm h)$-zero-sum $\sum_{i}c_{i}u^{(i)}=0$, where the $c_{i}$’s are not all zero. Since each $u^{(i)}$ is $(\pm h\to\pm\lfloor h/2\rfloor)$-reducible, we can replace $c_{i}u^{(i)}$ with a $(\pm\lfloor h/2\rfloor)$-sum of vectors from $S_{i}$. After the substitution, we obtain a $(\pm\lfloor h/2\rfloor)$-zero-sum. To see it is nontrivial, we recall that some $c_{i}$ is nonzero and the corresponding $u^{(i)}$ is also nonzero, which means $c_{i}u^{(i)}$ is nonzero and hence its substitution cannot be a trivial sum.
@@ 3.1 Simple $\mathrm{SIS}^{\infty}$ algorithms
We also remark that Theorem 3.9 (and even its weaker version Theorem 3.5) improves the result of [17] by a factor of $q^{O(q\log q)}$.
@@
$\sum_{i\in S}u_{i}^{\prime}-\sum_{i\in S^{\prime}}u_{i}^{\prime}=0,$
where $S,S^{\prime}\subseteq[m]$ are disjoint and not both empty. With Equation 11, we obtain the following zero-sum
$\sum_{i\in S}u_{i}^{\prime}+\sum_{i\in S^{\prime}}u_{i}+\sum_{i\in S^{\prime}}u_{i}^{\prime}=\left(\sum_{i\in S}u_{i}^{\prime}-\sum_{i\in S^{\prime}}u_{i}^{\prime}\right)+\sum_{i\in S^{\prime}}(u_{i}+2\cdot u_{i}^{\prime})=0,$
@@
This means $\sum_{i\in S_{j}}\alpha_{i}v_{i}=\beta_{j}v^{*}$ where $\beta_{j}=\sum_{i\in S_{j}}\alpha_{i}c_{i}$. Suppose there is some set $T\subseteq[d]$ that $\sum_{j\in T}\beta_{j}=1$. We can use the textbook dynamic programming algorithm to find such a set in time $\operatorname{poly}(q,d)$, then obtain a desired $(\pm 1)$-sum
@@
The reduction in [17] produces a bound of $m\geq q^{O(q\log^{2}q)}n^{O(q\log q)}$, whereas our Theorem 3.12 produces $m\geq n^{O(q)}$.
@@ 3.3 Simple $\mathrm{CIS}$ algorithms: large primes
Let $q\geq 3$ be a prime and let $A\subseteq\mathbb{F}_{q}$ be nonempty. Suppose there is a deterministic algorithm $\mathcal{A}$, given $m$ input vectors in $\mathbb{F}_{q}^{n}$ and running in time $T$, that outputs a nontrivial $A$-zero-sum. [^10] Since $\mathcal{A}$ works in the worst case and outputs nontrivial $A$-zero-sum, this immediately forces $0\in A$. Hence we do not need to explicitly assume $0\in A$ in the above statement.
@@
Firstly, a direct brute-force algorithm finds the AP in $A$ in time $\operatorname{poly}(q)$. Then we can assume without loss of generality that $A$ is the AP. By Lemma 3.13, it suffices to have a worst-case algorithm for nontrivial $(\pm\lfloor q/(2k)\rfloor)$-zero-sums, for which we use Theorem 3.9. ∎
In the case of large $q$, we have known results in arithmetic combinatorics that guarantee long APs. The following fact is due to Lev [25] and, in Appendix A, we provide a short proof for completeness.
Fact 3.16 ([25]).
@@ 4.1 Basic zero-sum algorithms for specific nontriviality
Denote $A=\{a_{1},\ldots,a_{c}\}$ where $c=|A|$. Let $\ell\leq n$ be the rank of the input vectors. By an invertible linear transform and in $\operatorname{poly}(n,r,|A|)$ time, we can assume that we are given standard basis $e_{1},\ldots,e_{\ell}$ and $v_{1}^{(i)},\ldots,v_{c}^{(i)},i\in[r]$ as $rc$ extra vectors that lie in the span of $e_{1},\ldots,e_{\ell}$.
@@
Since $\beta_{1},\ldots,\beta_{r}$ are not all zero, we can safely assume $\beta_{j}=1$ for some $j\in[r]$. Hence by substituting Equation 12 into Equation 13, we know that $a_{1},\ldots,a_{c}$ all appear as coefficients of $u_{\beta}$, which implies that $u_{\beta}$ is $A$-nontrivial.